OpenStudy (anonymous):

: What is the Area of a circle with a radius of 5. Leave pi in your answer.

OpenStudy (zzr0ck3r):

\(A= \pi*r^2\)

OpenStudy (zzr0ck3r):

so \(\pi*5^2 = \pi*25 = 25\pi\)

OpenStudy (anonymous):

so thats the answer

OpenStudy (anonymous):

@zzr0ck3r

OpenStudy (zzr0ck3r):

area = pi * r * r for you r = 5 so area = pi*5*5

OpenStudy (anonymous):

ok thx